OpenStudy (anonymous):

if 8+1=79, then 5+2+?

OpenStudy (anonymous):

the answer is 37

OpenStudy (saifoo.khan):

50

OpenStudy (saifoo.khan):

i guess.

OpenStudy (saifoo.khan):

37

OpenStudy (anonymous):

37??? makes sense

OpenStudy (anonymous):

the answer is 37.... 8+1=79 bcoz 8-1=7 and 8+1=9..... so in 5+2? 5-2=3 and 5+2=7....making the answer to be 37......

OpenStudy (saifoo.khan):

:D

OpenStudy (anonymous):

37

jhonyy9 (jhonyy9):

5+2+2 = 79